eviant:

Determine the general form of the equation for the circle x^2 + (y + 1)^2 = 2.

563blackghost:

The formula for a circle is: \(\large\bf{(x-h)^{2}+(y-k)^{2}=r^{2}}\) \(\bf{(h,k)}\) is your vertex. \(\bf{r}\) is your radius. So determine from there.

eviant:

x2 + y2 + y – 1 = 0

eviant:

@dude @Vocaloid @TheSmartOne @563blackghost

sillybilly123:

*general form* is written as \(f(x,y) = 0\) so multiply/ie FOIL it all out and set LHS to zero

eviant:

@Shadow is my answer correct?

Shadow:

\[x^2 + (y + 1)^2 = 2\] \[x^2 + (y + 1)(y + 1) = 2\] \[x^2 + y^2 + 2y + 1 = 2\] \[x^2 + y^2 + 2y - 1 = 0\]

Shadow:

\[1 \times y = y\] You do this 2x, and them add them \[y + y = 2y\]

eviant:

so, the answer is x^2+y^2+2y-1=0?

Shadow:

Yes. Do you see how I got there?

eviant:

yes, thanks for explaining it so well

eviant:

though, how do you get 2y, in the third step? is it blc there are two Ys?

eviant:

@Shadow ?

Shadow:

FOIL means First, Outer, Inner, and Last We have: \[(y + 1)^2 \] This is also represented as: \[(y + 1)(y + 1)\] We multiply the first: \[y \times y = y^2\] We multiply the outer: \[y \times 1 = y\] We multiply the inner: \[1 \times y = y\] We multiply the last: \[1 \times 1 = 1\] Sum them, or add them up: \[y^2 + y + y + 1 = y^2 + 2y + 1\]